A monochromatic source of light operating at 200 W emits 4 x 1020 photons per second. Find the wavelength of the light.

Grade:12

1 Answers

Arun
6 years ago
The energy of each photon=
E = (200 J/sec)/(4 * 10^20 /sec)
= 5 * 10^-19 J
 
Now \lambda = hc/ E
\lambda = (6.6* 10^-34) *(3* 10^8)/(5 * 10^-19)
= 400 nano metre